More Exercises 1. What mass of carbon dioxide can be made from 6g of carbon? (equation C +02-CO2) 2. What mass of sulphuric acid can be made from 8g of SO (equation SO3 + Hz。= H2SO4) 3. What mass of HCI reacts with 6g of magnesium? (equation Mg+ 2HCI-MgC2+H2) 4. What mass of water is made when 32g of methane burns? (equation CH4+202 CO2+ 2H2O) 5. What mass of hydrogen reacts with 14g of nitrogen? (equation N2+3H2 2NH) 6. What mass of Sulphuric acid reacts with 80g of NaOH? (equation H SO4+ 2NaOH-Na SO4+2H2O) 7. What mass of propane will react with 80g of oxygen? (C,Hs+502 3C02+4H20) Hi, I need help with the answers for #1-7
Answer #1

1. C+O2 --> CO2

1 mol C = 1 mol CO2

NO of mol of C reacted = 6/12 = 0.5 mol

   NO of mol of Co2 FORMED = 0.5 mol

amount of CO2 formed = 0.5*44 = 22 g

2) so3 + H2O ---> H2SO3

   1 mol SO3 = 1 mol H2SO3

No of mol of SO3 = 8/80 = 0.1 mol

no of mol of H2SO3 FORMED = 0.1

amount of h2so3 = 0.1 *82 = 8.2 g

3) 1 mol Mg = 2 mol HCl

no of mol of Mg = 6/24 = 0.25 mol

no of mol of HCl = 0.25*2 = 0.5 mol

mass of hCl reacted = 0.5 *36.5 = 18.25 g

4) 1 mol CH4= 2 mol h2o

no of mol of CH4 = 32/16 =2 mol

amount of water = 2*2*18 = 72 g

5) 1 mol N2 = 3 mol H2

no of mol of N2 = 14/28 = 0.5 mol

mass of H2 reacts = 0.5*3*2 = 3 g

6) 1 mol H2SO4 = 2 mol NaOH

   NO OF mol NAoh = 80/40 = 2 MOL

   no of mol of H2SO4 = 2/2 = 1 mol

mass of H2SO4 = 1*98 = 98 g

7) 1 MOL C3H8 = 5 mol O2

   no of mol of O2 = 80/32 = 2.5 mol

   mass of propane = 2.5/5*44 = 22 g